A pebble is shot vertically at 28m/s ignoring air resistance. What height will the pebble reach?


1
Expert's answer
2020-02-10T09:08:02-0500

From the conservation of energy:


"0.5mv^2=mgh"

"h=\frac{v^2}{2g}=\frac{28^2}{2(9.8)}=40\ m"